OpenStudy (anonymous):

Violet was gathering blueberries along a path from her back door to the rear of her garden. On the way towards the rear of her garden, she walked at an average speed of 6 feet per minute and on the way back to her back door, she walked at an average speed of 15 feet per minute. If the round trip took 28 minutes, how far in feet is it from her back door to the rear of her garden?

OpenStudy (anonymous):

let x be the distance of her house from the rear garden, then we hav x/6+x/15=28...solve for x,u get distance in feet.

OpenStudy (anonymous):

x=120 feet

OpenStudy (anonymous):

6x-15y=0 (0 since she is right back where she is ) x+y=28

OpenStudy (anonymous):

I averaged the total speed, then multiplied by 28...

OpenStudy (anonymous):

so which way is right?

OpenStudy (anonymous):

(6+15)/2=10.5 feet/min average speed 10.5x24=294

OpenStudy (anonymous):

6(28-y)-15y=0 168-6y-15y=0 168=21y y=8 minutes Distances= time * speed 8 * (15)=120

OpenStudy (anonymous):

d=r*t

OpenStudy (anonymous):

okay you might be right...

OpenStudy (anonymous):

nope u should not average the speed and take it,for distance of x. u should proceed the way i've given...for correct answer.